Find the equation of the line passing through the point (-1,-4) that is parallel to the line -9+5y=6
So far I have 5y=-9+6 but I am not sure because there is no x.

Answer by Alan3354(69443)

Find the equation of the line passing through the point (-1,-4) that is parallel to the line -9+5y=6
So far I have 5y=-9+6 but I am not sure because there is no x.
----------------
Sometimes there is no x.
5y=-9+6
5y = -3
y = -3/5
---
That's a line parallel to the x-axis thru (-3/5,0)
Its slope is zero.
All parallel lines have the same slope.
----
A line thru (-1,-4) with a slope of zero is y = -4
It's parallel to the x-axis.
y = -4 for all values of x.